# Towards a multi-dimensional picture of nucleation: The example of liquid condensation through theory and rare event sampling

### v3

@@ Abstract
Conclusions - In a context where CNT is being increasingly challenged by experimental and simulation observations, we revisited the longstanding yet still relevant problem of liquid condensation. To begin, we propose innovative simulation approaches to probe the nucleation event and especially collect structures that are precisely situated at the saddle point of the free energy barrier. In order to span a large density range, we considered both an original implementation of the seeding approach and a combination of steering MD with aimless shooting. Altogether, our simulation results show that during liquid condensation, the density of the nucleating clusters differs from that of the equilibrium liquid and decreases with the increase of vapor density. Then, we expanded the capillary model to include the density of the nucleating cluster as a supplementary driving parameter. Our straightforward two-variable nucleation theory is self-consistent since all of its parameters are not fitted to any nucleation simulations and are separately obtained through physical considerations. Remarkably, our model successfully captures the concomitant growth and densification of nearly critical cluster, as observed in our simulations, and also achieves quantitative agreement for nucleation rates and critical cluster densities. As such, by providing highly accurate predictions without adjustable parameters, our two-variable model offers a powerful alternative to CNT, with broad applicability to diverse nucleation processes.
@@ Appendix A Simulation protocol
Simulations were performed using the open-source package LAMMPS [88]. For the Lennard-Jones interactions, we used a large cutoff of $6.78\sigma$, ensuring excellent agreement between the phase diagram obtained from perturbation theory and that from simulations [89, 74]. The employed timestep is $0.005\tau$, with $\tau=\sigma m^{\frac{1}{2}}\epsilon^{-\frac{1}{2}}$. Specifically, we work at $T=0.8\epsilon/k_{\mathrm{B}}$. For temperature and pressure controls, we used the Nose-Hoover thermostat/barostat implemented in LAMMPS with damping parameters taken respectively as $0.5\tau$ and $5\tau$.
@@
At low supersaturation, the NVT seeding technique is employed to stabilize liquid clusters in equilibrium with the vapor [44]. Once the cluster is stabilized, the simulation is converted to the NPT ensemble to emulate an infinite system. In this stage, the pressure is prescribed as the time-averaged virial pressure from the preceding NVT simulation [90]. The system volume becomes a free variable, accommodating itself to maintain the pressure. As the considered cluster is critical (but now unstable), it has an equal probability ($\sim 0.5$) of growing, from which growth rates can be estimated, or shrinking. To identify critical clusters and their growth rates, multiple NVT-to-NPT runs are performed from different initial configurations, varying particle positions and velocities while keeping the desired quantities. In the NVT ensemble, we were able to stabilize 20 liquid clusters per box size ($20\sigma$, $50\sigma$ and $90\sigma$). When checking their criticality in infinite systems, 20 NPT runs were performed.
@@
In intermediate supersaturation regimes, we used a set of biasing techniques. Here are some details on the implementation. Firstly, for the steered MD, simulations were run for $10^{4}\tau$ with $10^{4}$ atoms. For the collective variable, we used $N_{\mathrm{liq}}$ defined as the number of atoms having coordination number larger than 6. The constraint potential constant was chosen as $100/s^{2}$ where $s$ is the standard deviation of $N_{\mathrm{liq}}$ measured in the vapor phase. Steered MD were performed three times for each of the considered vapor density to ensure repeatability. Secondly, from the steered MD accumulated structures, we measured the commitment probability to obtain brute-force trajectory connecting vapor and liquid phases using 10 different initial velocity conditions. For each of the three steered MD simulations, we obtained one connecting trajectory. Thirdly, the obtained trajectories were used as initialization for three-point aimless shooting algorithm [54] where we sampled configurations separated by a time interval equal to $10\tau$. From each of the connecting trajectory, we launched 10 different aimless shooting sampling. Finally, we measured the commitment probability of every structures obtained with aimless shooting using 50 initial velocities. Structures were considered as critical when there commitment probability is located in the range $[0.3,0.7]$. More details can be found in Ref. [91].
@@ Appendix B Simulation analysis
Cluster radius $R$, density $\rho$, and vapor density $\rho_{0}$ from simulations were determined by fitting the integral of a modified sigmoid function [92],
@@
where $A$ represents the liquid/vapor interface width and $b=2A/(\exp{(2AR)}+1)$, to the integrated radial density profile calculated from the center of the stable cluster. Post-processing was implemented in Python, utilizing Ovito’s MD analysis library [93].
@@ Appendix D Classical nucleation theory (CNT) predictions
Finally, the nucleation rate predicted by CNT is given by [56, 55]
@@ Appendix E Incorporation of dynamical density functional theory (DDFT) into nucleation theory
In DDFT, the driving force for particle diffusion is assumed to be proportional to the gradient of the chemical potential in the real space $(\mathbf{r})$ [63],
@@
Following closely the method proposed by Lutsko [62], we parameterize the density profile in order to derive a dynamical equation in the phase space $\{\boldsymbol{x}\}$ for the nucleation theory. We define $n(r,t)$ as the integrated density within a spherical shell of radius $r$,
@@
This is the $\mathbf{g}$ matrix introduced by Lutsko [62] in its so-called mesoscopic theory of nucleation. $\mathbf{g}$ can be directly connected to Alekseechkin’s theory [56]. The right-hand side of Eq. (8) can be evaluated using integration by parts, yielding
@@
we can directly identify the diffusivity matrix as $\mathbf{D}=D_{\mathrm{diff}}\,\mathbf{g^{\mathrm{-1}}}$. For the two-variable model, the diffusivity matrix is given by [59]:

### v4

@@ Abstract
Conclusions - We revisited the longstanding yet still fundamental problem of liquid condensation and uncovered a non-classical nucleation pathway. Using innovative rare-event simulation strategies – an original implementation of the seeding approach together with steering MD combined with aimless shooting – we were able to probe nucleation events with unprecedented precision and collect structures situated at the saddle point of the free energy barrier across a wide range of supersaturation. Our simulation results show that the density of the critical clusters differs markedly from that of the equilibrium liquid and decreases as the vapor density increases. Building on this key insight, we expanded the capillary model by explicitly incorporating the cluster density as a supplementary driving parameter. Without any fitted parameters, this self-consistent two-variable nucleation theory successfully captures the concomitant growth and densification of nearly critical cluster, as observed in our simulations, and also achieves quantitative agreement for nucleation rates and critical cluster densities. Taken together, these results demonstrate that cluster density is an essential collective variable governing nucleation, and that omitting it is a principal reason for the longstanding failures of CNT. By establishing a multidimensional picture of nucleation, our work provides a broadly applicable framework for refining nucleation theories.
@@ Appendix A Simulation protocol
Simulations were performed using the open-source package LAMMPS [LAMMPS]. For the Lennard-Jones interactions, we used a large cutoff of $6.78\sigma$, ensuring excellent agreement between the phase diagram obtained from perturbation theory and that from simulations [baidakov_self-diffusion_2011, baidakov_metastable_2007]. The employed timestep is $0.005\tau$, with $\tau=\sigma m^{\frac{1}{2}}\epsilon^{-\frac{1}{2}}$. Specifically, we work at $T=0.8\epsilon/k_{\mathrm{B}}$. For temperature and pressure controls, we used the Nose-Hoover thermostat/barostat implemented in LAMMPS with damping parameters taken respectively as $0.5\tau$ and $5\tau$.
@@
At low supersaturation, the NVT seeding technique is employed to stabilize liquid clusters in equilibrium with the vapor [Vega-Seeding-NVT]. Once the cluster is stabilized, the simulation is converted to the NPT ensemble to emulate an infinite system. In this stage, the pressure is prescribed as the time-averaged virial pressure from the preceding NVT simulation [PhysRevE.101.022611]. The system volume becomes a free variable, accommodating itself to maintain the pressure. As the considered cluster is critical (but now unstable), it has an equal probability ($\sim 0.5$) of growing, from which growth rates can be estimated, or shrinking. To identify critical clusters and their growth rates, multiple NVT-to-NPT runs are performed from different initial configurations, varying particle positions and velocities while keeping the desired quantities. In the NVT ensemble, we were able to stabilize 20 liquid clusters per box size ($20\sigma$, $50\sigma$ and $90\sigma$). When checking their criticality in infinite systems, 20 NPT runs were performed.
@@
In intermediate supersaturation regimes, we used a set of biasing techniques. Here are some details on the implementation. Firstly, for the steered MD, simulations were run for $10^{4}\tau$ with $10^{4}$ atoms. For the collective variable, we used $N_{\mathrm{liq}}$ defined as the number of atoms having coordination number larger than 6. The constraint potential constant was chosen as $100/s^{2}$ where $s$ is the standard deviation of $N_{\mathrm{liq}}$ measured in the vapor phase. Steered MD were performed three times for each of the considered vapor density to ensure repeatability. Secondly, from the steered MD accumulated structures, we measured the commitment probability to obtain brute-force trajectory connecting vapor and liquid phases using 10 different initial velocity conditions. For each of the three steered MD simulations, we obtained one connecting trajectory. Thirdly, the obtained trajectories were used as initialization for three-point aimless shooting algorithm [Peters2006Aug] where we sampled configurations separated by a time interval equal to $10\tau$. From each of the connecting trajectory, we launched 10 different aimless shooting sampling. Finally, we measured the commitment probability of every structures obtained with aimless shooting using 50 initial velocities. Structures were considered as critical when there commitment probability is located in the range $[0.3,0.7]$. More details can be found in Ref. [Lam2024].
@@ Appendix B Simulation analysis
Cluster radius $R$, density $\rho$, and vapor density $\rho_{0}$ from simulations were determined by fitting the integral of a modified sigmoid function [Lutsko_2008],
@@
where $A$ represents the liquid/vapor interface width and $b=2A/(\exp{(2AR)}+1)$, to the integrated radial density profile calculated from the center of the stable cluster. Post-processing was implemented in Python, utilizing Ovito’s MD analysis library [ovito].
@@ Appendix D Classical nucleation theory (CNT) predictions
Finally, the nucleation rate predicted by CNT is given by [alekseechkin_multivariable_2006, langer_statistical_1969]
@@ Appendix E Incorporation of dynamical density functional theory (DDFT) into nucleation theory
In DDFT, the driving force for particle diffusion is assumed to be proportional to the gradient of the chemical potential in the real space $(\mathbf{r})$ [Evans01041979],
@@
Following closely the method proposed by Lutsko [lutsko_communication_2011], we parameterize the density profile in order to derive a dynamical equation in the phase space $\{\boldsymbol{x}\}$ for the nucleation theory. We define $n(r,t)$ as the integrated density within a spherical shell of radius $r$,
@@
This is the $\mathbf{g}$ matrix introduced by Lutsko [lutsko_communication_2011] in its so-called mesoscopic theory of nucleation. $\mathbf{g}$ can be directly connected to Alekseechkin’s theory [alekseechkin_multivariable_2006]. The right-hand side of Eq. (8) can be evaluated using integration by parts, yielding
@@
we can directly identify the diffusivity matrix as $\mathbf{D}=D_{\mathrm{diff}}\,\mathbf{g^{\mathrm{-1}}}$. For the two-variable model, the diffusivity matrix is given by [lutsko_two-parameter_2015]: